OpenStudy (anonymous):

graph this equation: y - 2 = 2/3(x - 4)

OpenStudy (mathteacher1729):

Is that 2/3 times (x-4) or 2 over (3*(x-4)) all in the denom?

OpenStudy (anonymous):

no, the equation is y - 3 = 2/3 times x-4

OpenStudy (anonymous):

so just 2/3 times (x-4)

OpenStudy (mathteacher1729):

Ah, ok. So here is the SUPER DUPER important fact about graphing lines: You need only TWO points to make a line. (y - 2) = 2/3(x - 4) So when x = +4 that means y -2 = 0 (or, y = 2) POINT ONE: (4,2) The SLOPE of this line is 2/3. That means we go "up 2 and over 3" from any point on the line. let's go up 2 and over 3 ... starting from (4,2) ... that leaves us at (7, 4). POINT TWO: (7,4). Now just plot those two points and draw a line between them. :) Let me know if you have any questions about the process.

OpenStudy (anonymous):

thank you so much! thats perfect i get it, what about y + 4 = -3(x + 2)?